OpenStudy (anonymous):

Factor completely 2x^5+ 10x^4– 22x^3 Factor completely 2x^5+ 10x^4– 22x^3 @Mathematics

OpenStudy (hoblos):

2x^3 (x^2 +5x -11)

OpenStudy (agreene):

\[\frac{-1}2 (-2 x+\sqrt{69}-5) x^3 (2 x+\sqrt{69}+5)\] is the lowest possible factorization.